OpenStudy (anonymous):

What is the probability that, in an organism with a haploid number of 12, a sperm will be formed that contains all 12 chromosomes whose centromeres were derived from maternal homologs?

OpenStudy (joppei):

Most sperm cells in the animal kingdom are haploid so lets say that the normal organism is diploid (has 2*12= 24 chromosomes). When a sperm cell is formed, one chromosome of each chromosome pair is put into the sperm cell. Which makes it haploid and thus gives it 12 chromosomes. If you have 2 chromosomes which are divided, the chance of getting one of them is 50%. Now, if you have 12*2 chromosomes which are divided this means that if you want to get the fathers chromosome every time this 50% chance must have the same outcome 12 times in a row. This gives you the formula: 50%^12 (50%*50%*50%*50%*50%*50%*50%*50%*50%*50%*50%*50%*) or as it is easily calculated: 0.5^12 (In which 1 states it happens every time and 0.5 states it happens half of the times etc.) So the answer is; 0.5^12=2.4*10^-4 or: 0.00024 in other words 1/0.00024 gives us a chance of 1 in 4096